OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and third equations, then the first and second equations. 4x + y + 3z = 12 5x – 2y – 2z = –11 –x – y – 6z = –9 a. 3x – 3z = 3 13x + 4z = 13 b. 3x – 3z = 3 x + 6z = 9 c. 4x + 3z = 12 x + 6z = 9 d. 4x + 3z = 12 13x + 4z = 13

OpenStudy (anonymous):

add 1st and 3rd: gives 3x - 3z = 3 1st and 2nd multiply 1st by 2 ; 8x + 2y + 6z = 24 5x - 2y - 2z = -11 add 13x + 4z = 13 option a is correct

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

yw